Yesterday Tween Teen Fashions had seven less than three-fourths of its sales transactions paid for by credit cards. If 209 transactions were charged, how many total transactions took place? Would it be 7-3/4x= total transactions?? Really confused on this one
Answer by srinivas.g(540)

x= total no of transactions
7 less than 3/4 th transactions means 3/4 x -7
which ie equal to 209 ( charged transactions)
so 3/4 x - 7 = 209
3/4 x = 209+ 7
3/4 x = 209 + 7
3/4 x = 216
x = 216 *4/3
=288
total transactions = 288